A particle starting from rest moves with angular velocity omega =a-bt. Find the total angular swept by the particle before coming to rest

Grade:11

1 Answers

abhinav agarwal
6 years ago
ω= a-btwhen ω=0therfore t=a/bdθ/dt=a-btdθ=(a-bt)dtintegrating∫dθ=∫(a-bt)dt from 0-a/bwe getθ=a^2/2b